OpenStudy (anonymous):

Find the derivative with and without using the chain rule. f(x) = (x^2 + 1)^3

OpenStudy (anonymous):

\[f(x)=(x^2+1)^3=(x^2+1)(x^2+1)(x^2+1)\] \[f'(x)=6x(x^2+1)^2\]

OpenStudy (anonymous):

used only product rule

OpenStudy (anonymous):

is that using the chain rule?

OpenStudy (anonymous):

no. i said i only used product rule.

OpenStudy (anonymous):

okay

zepdrix (zepdrix):

One, do you understand how to find the derivative `with` using chain rule? You would just take the derivative as is.

OpenStudy (anonymous):

not 100%

OpenStudy (anonymous):

okay i'll try doing with the chain rule

zepdrix (zepdrix):

Ok :)

OpenStudy (anonymous):

okay can i post it here and correct me if i i mistake somewhere

zepdrix (zepdrix):

k

OpenStudy (anonymous):

\[d/dx ((x^2 + 1)^3) = 3(x^2 + 1)^2 d/dx(x^2 + 1) = 3(d/dx(x^2)(x^2 - 1)^2 = 3(2x)(x^2 = 1)^2 = 6x(x^2 + 1)^2 \]

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

I'm not quite sure why it changed from x^2+1 to x^2-1. Was that a typo?

OpenStudy (anonymous):

i see that it didn't show up all of it but here is the last part 3(2x)(x^2 + 1)^2 = 6x(x^2 + 1)^2

OpenStudy (anonymous):

yes typo for that one

zepdrix (zepdrix):

\[\large f'(x)=3(x^2+1)^2\frac{d}{dx}(x^2+1)\]\[\large f'(x)=3(x^2+1)^2(2x) \qquad \qquad f'(x)=6x(x^2+1)^2\] Ok yah looks good :)

OpenStudy (anonymous):

thanks!

zepdrix (zepdrix):

`without` chain rule is going to be a little tricky. They want you multiply out your binomial :O oh boy.

OpenStudy (anonymous):

ohh so the one @ingenuus showed is nto by without chain rule?

OpenStudy (anonymous):

not*

OpenStudy (anonymous):

the one ingenuus showed its using the product rule

zepdrix (zepdrix):

That's another method for doing it without chain rule. Which involved the product rule with `3 terms`. Which is a little bit tricky if you've never seen it before. Product rule with 2 terms is probably what you're used to seeing. I was referring to a different method, multiplying out all the brackets as Wio showed you the last time you posted this. :D

OpenStudy (anonymous):

Ohhhh okay got you